Fermi National Laboratory
 

Accelerator Update for August 17, 2007 - August 24, 2007

 

Fermilab Shutdown Report

8/17/07 – 8/24/07


LINAC

  • All work on Schedule

  • LLRF Upgrade

Booster

Electrical Work:

  • Cable pull 1000% complete

Water work:

  • LCW water work in the Booster West Gallery
    • Area continues to be a working area
    • 6” copper pipe successfully pressure tested and mounted on the wall
    • 4” copper pipe and RF manifold upgrade progressing well
    • Next week: Continue with the installation
  • Tunnel water work:
    • 6 filter canisters installed in the tunnel (L2, L4, L11, L14, L16, L22)
    • Plans continue to be developed for Booster de-oxy system
    • Next week: Connect up water to the correctors
      • Continue with the filtration system installation (piping)
  • CUB
    • VFD test occurred most of the week (1 pump only)
  • Corrector installation
    • Number of new correctors in the tunnel: 10
      • L19, L17, L21, L23, L15, L9, L11, L13, L5, L7
      • They are all fully surveyed and vacuum connected
    • Booster vacuum slowly recovering
    • Next week:
      • Continue with corrector installation
      • Continue monitoring vacuum
      • Preparatory work needs to be completed for the remainder locations (critical ones…L1 and L3).
  • Additional Activities
    • FESS
      • Burned out lights were replaced in the tunnel
      • Sump pump at L2 was replaced
      • Engineers evaluate tunnel structure
        • Cracks developed in the ceiling at S5 and P16 alcove
      • Sinkhole is still a sink hole…
        • Lots of effort to understand the discharge pattern…
  • Power Outage (Feeder 41) schedule for Aug 25 th
    • Booster will be switched to back feed from CUB this afternoon.
    • Sumps and lights will be on during the outage.
  • RF maintenance work started.

Antiproton Source

Mechanical

  • Leak Check (4)
    • Finished checking D/A, A40, A30
      • D:ESEP, A30 IP fixed
  • D:ISEP Shielding
    • Stand set; started stacking Deb DCCT motorized stand
  • Ion Chamber removal
  • Deb DCCT motorized stand
  • D:EKIK ceramic tube
  • AP2 Water Hose Replacement
  • Leaking LQA
    • New one in beam line
  • Replaced leaking Deb Sextuple
  • SMA water leak
    • As Found Survey
  • 4 water leaks to fix
    • IQ8 fixed twice
  • Cryo
    • Valve Stem replace
    • Leak Check (1 of 2)
    • LN2 transfer line maintenance

Target Station

  • Water Systems (2)
    • DI cartridge replaced
  • Temperature Monitoring (2)
  • Target Rotation
    • Removed from vault; see photo
  • Storage
  • SEM Module
    • Removed from vault; see photo
  • LVDTs
  • Sample Retrieval
  • Lens motion
  • Air Exhaust Stack
  • 10mm-2 into transformer
    • Prepping; progress
  • 10mm-3 into transformer
  • prepping
  • Lens replacement
    • Pictures of current lens
  • Pictures of PMAG

Cooling Systems

  • Deb 3&4 Trans Notch Filters
  • Deb Double Notch Filters
  • Core 4-8GHz Equalizer
  • TWT balance & trip points
  • Stacktail Filter 3 cable change

Electrical (Pbar & EE Support)

  • Shunt Move
    • Started
  • Shunt Repairs (6)
    • Removed and rebuilding
  • PS Maintenance
  • Lens SCRs
  • Regulation Electronics
  • Ion Pump maintenance
  • Deb Shunt re-arrangement
    • Plan has been proposed

Other

  • Interlocks/Safety System (3)
  • Repair cathodic protection
  • Dirt holes under buildings
    • Dripping from tunnel ceiling
  • DCCT calibration
  • Baking System checkout
  • RF & TWT filters maintenance
  • Lights Checkout
    • FESS to fix
  • Acc BPM repairs
  • Activation Survey
    • Finished AP1, Acc, Deb, AP2
  • Inductance checkout
  • PM ODH Fans & Sumps
  • Tunnel Wireless
  • Install Leak Detector
  • SEM maintenance
  • Clogged sump line – Transport
    • Requires berm dig

Main Injector

Pictures of collimator build

Miscellaneous maintenance

Recycler

  • Finished or underway
    • New DCCT section installed, leak checked and under vacuum since Wednesday last week; pumping down.
    • DCCT section has been aligned
    • V303 moved, need to make connection and complete shielding
    • H308 moved; need to make connection and complete shielding
    • Need to check ALL LEP dipole positioning wrt. beampipe
    • STC tank 7 leak check underway
    • Trombone additional quadrupoles installation issues
      • HLRF cooperation
      • Finished: 601,602,606,607
      • Stands ready: 603,604,608
      • Moved MI buss support at 605
      • Wait on 609 location (radiation)
    • Added support to old DCCT
    • Bulk supply upgrade (on-going)
    • DCCT cable connected
  • On the list
    • Magnet moves (305, 307) and survey alignment
    • Elevation survey before and after collimator installation
    • 800 transfer line vacuum work
    • Vacuum maintenance
    • Kicker repair; timing verification
    • TOR330 repair
    • BPM calibration
    • Diagnostics maintenance
    • MI-21 SA camera/TV repair
    • Laser link vacuum maintenance
    • Additional quad cabling and field check

Tevatron

Tevatron status as of August 24

SwitchYard

Continuing with maintenance work.

NuMI

  • Completed water sampling August 20th
  • Drain cleaning presently taking place in the pre target
  • Installed new target pile air system humidity sensors
  • Changed Pre & HEPA filters in target pile Air System
  • Installed protective roof covers on strip line cage
  • Re-started Horn 2 LCW and started leak investigation
  • Target Pile Dehumidification Installation Project
    • Completed installation of PVC condensate line in shaft
    • Completed installation of electrical conduit in shaft
    • Continue with chilled water pipe installation @ bottom of shaft
    • Completed layout of 4” chilled water pipe in Upper Labyrinth
    • Started installation of 4” chilled water pipe hangers in Target Hall
    • Started planning & layout for duct work to Munter dehumidifiers

MiniBooNE

There was no work completed last week.

  • Reverse Horn Polarity
    • Take wipes.
    • Reverse polarity of horn power supply.
    • Test horn.
  • Check Corrosion Tabs at 25m Absorber Next Week
    • Unstack 9 blocks from 25m absorber.
    • Retrieve corrosion tabs.
    • Restack blocks.
    • Seal cover.
  • Change RAW DI Bottle (50%)
  • Dry out horn (09Aug07)
  • Check Leakage Current on Ion Pumps
  • Power Supply Maintenance
  • Toroid Work

MTA Beamline Installation

  • Tunnel activities
    • Janitors cleaned up the garage area
    • Contractors started working on the cable tray installation
    • Moved an existing turbo cart out of the way
      • Some beam line components have to be installed in that area
  • Shielding underneath hatch area
    • Crane showed up last Friday (Aug 17 th)
      • Shielding blocks were completely removed by early Monday morning (1.5 days)
      • Crane was heavily used for:
        • Lift 3 barrels full of concrete (old ramp)
        • Lowering auxiliary equipment to aid the magnet installation
        • Lowering all magnets in the tunnel
        • Lowering all the LCW pipes, cable trays
        • Lowering most of the big cable spools
  • Next week:
    • Crane/Iron Workers will be moved to another job site.
      • Expect to return Tuesday (08/28)
      • Parking lot will remain roped off during this time.
    • Leveling the floor will occur during this time.
  • Electrical Work

  • Alignment
    • Completed the network and stake out the MTA stub area
    • Assisted on measurements for the core drill
    • Next week:
      • Check magnets stationing and elevation
  • Core Drill
    • Core drill of 3 blocks was successful performed yesterday
      • Some other cuts were also done in order to get the right amount of shielding blocks for the new wall.
  • Magnet Installation
    • Being performed by Iron Workers
      • 40% magnets in place

        • Remaining will be stationed today
    • Next week: proceed with spool pieces after elements have been surveyed.
  • LCW work
    • Pipe fitters start their activities today!
      • 1 day ahead of schedule! Not too bad for the MTA first week of tunnel activities.

Collider Detector at Fermilab

  • Plugs bores sealed, dewpoint at -10 °C.
  • Open up silicon to the outside with it cold
  • ISL cooling repair commenced yesterday
  • Supply and return lines cut and straightened
  • Short straight shot to repair area with borescope and glue tube
  • Storms: emergency radio didn’t seem to work, some flooding in the building

D-Zero

Toroid water leak repaired.

Alignment

Completed Jobs

  • Recycler Task Manager: Oshinowo
    • 5527 Optically align Recycler DCCT after installation. Date: 8/13/2007 Complete
  • Tevatron Task Manager: Bocean
    • 5494 As found low beta quads and separators at B0 Date: 8/06/2007- 8/08/2007 Complete
    • 5495 As found low beta quads and separators at D0. Date: 8/09/2007- 8/14/2007 Complete
    • 5498 Measure rolls on 111 dipoles and quads in Tevatron (48) Complete
    • Un roll D16-1 quad when D-1 is warmed.(As Found) Date: 8/13/2007 Complete
    • 5591 As Founds of the Roman pots at D0 4 of 6 Date: 8/13/2007 85% Complete
    • A0 dump as found (no request for this job) Date: 8/23/2007 Complete
  • Booster Task Manager: Oshinowo
    • 5534 Align 24 Booster Corrector Magnets.( Long 19 Long 17 ) Date: 8/13/2007 Complete
    • 5534 Align 24 Booster Corrector Magnets ( Long 21 Long 23 ) Date: 8/14/2007 Complete
    • 5534 Align 24 Booster Corrector Magnets ( Long13 Long15 ) Date: 8/20/2007 Complete
    • 5534 Align 24 Booster Corrector Magnets ( Long 11 Long 9 ) Date: 8/21/2007 Complete
    • 5534 Align 24 Booster Corrector Magnets ( Long 7 Long 5 ) Date: 8/23/2007 Complete
  • Main Injector Task Manager: Oshinowo
    • 5563 AS found survey of Quadrupoles at 400, 402, 404 and the three Lambertson. Date: 8/13/2007- 8/15/2007 Complete
    • 5554 307, 308 Collimators: Base plate survey Date: 8/16/2007 Complete
    • 5554 307, 308 Collimators: Set travel and limit switches Date: 8/22/2007 Complete
    • 5596 Magnet 614 had to be moved please realign. Date: 8/23/2007 Complete
  • CMS Task Manager: Oshinowo Field Crew: O’Boyle.
    • 5516 Survey the CMS F-pixel Half cylinder +Z2 in the Lab C Date: 8/09/2007 Complete
  • D0 Task Manager: Oshinowo
    • Stickmike survey of Muon system prior to opening of detector. This will include both CFs and EFs. Date: 8/06/2007 Complete
    • 5524 VStar or Laser Tracker survey of North/South EF pedestals with respect to the sidewalk monuments. Date: 8/07/2007 Complete
    • 5592 Elevation run in D0 before machine is opened. Date: 8/06/2007 Complete
  • Linac Task Manager: Oshinowo
    • 5468 Stake out the MuCool beamline component locations in the Linac and MuCool Test Area (MTA). Date: 8/08/2007 Complete
  • P-Bar Task Manager: Oshinowo
    • A4Q10 (LQA in Accumulator) has a water leak. An "as found“ Date: 8/09/2007 Complete
    • 5595 An Accumulator multipole magnet (SMA) has developed an internal Date: 8/23/2007 Complete
  • Nova Task Manager: Oshinowo

Next week’s schedule of work

  • Tevatron Task Manager: Bocean Field Crew: Bradford, Adkins, J Wilson.
    • 5498 Measure rolls on 43 dipoles and quads in Tevatron. These are magnets that cannot be measured with roll fixture. Date: 8/14/2007- 8/27/2007
    • 5598 A0 dump as found (late request for this job) Date: 8/24/2007
    • 5591 As Founds of the Roman pots at D0 4 of 6. Date: 8/24/2007
    • 5506 As found crystal collimator at E0. Date: 8/27/2007
    • As found TEL1 in transfer hall. Date: 8/27/2007
    • As found TEL2 in transfer hall. Date: 8/27/2007
    • 5594 As found Schottkies, Separators A17 straight. Date: 8/24/2007 Manager: Kyle
    • Elevation run A-E sectors Complete Date: 8/14/2007- 8/28/2007
  • Booster Task Manager: Oshinowo
    • 5534 Align 24 Booster Corrector Magnets Date: 8/27/2007
  • Main Injector Task Manager: Oshinowo
    • 5554 4 collimators in the Main Injector at the 301, 303, 307, and 308 locations. Date: 8/24/2007-8/30/2007
  • Nova Task Manager: Oshinowo
    • 5475- Stake out new Recycler injection line (MI10)Date: 8/27/2007
    • 5477-Stake out new Recycler RR30 quad positions for the Nova ProjectDate: 8/29/2007
    • 5476-Stake out new Recycler extraction line (MI30) for the Nova Project Accelerator upgrade. Date: 8/30/2007
  • MTA Task Manager: Oshinowo
    • Shoot MTA shielding blocks after re-stacking them underneath the hatchway area.
    • 5469 Rough alignment of the MuCool beam line components. Date: 8/29/2007
  • CDF Task Manager: Bocean Not Scheduled
  • ILC Task Manager: Bocean Not Scheduled
  • P150 Task Manager: Oshinowo Late Request
    • 5593 We had to move several magnets and multiwires to make up the ends on the cans for vacuum connections. They are as follows: The multiwire cans at 708 and 710. B2 magnets at the interface of 710 - 711 and the interfaces of 709 - 710, the serial #'s are BDM 883-1 BDM 835-1 BDM 879-1 and BDM 397-1. Date: 8/28/2007

Alignment and Metrology Home page

Mechanical

Janitorial work in the tunnels continues.

Cryogenic System

Tevatron at LN2 temperature except C4, D1, and E1 at room temperature

  • Tunnel magnet cryostat vacuums at A15, A25, and D35 regions are of concern.
  • TeV Cryogenic Transfer Line F2-F3 vacuum leak checked fine.
  • CDF Cryo back on TeV Cryo Compr discharge/suction headers due to pond work

Cryo Operations

  • 3 FRIG Mycom compressor (plus DZero & CDF Cryo) running with helium precooled to LN2 temperature flowing through magnet strings.
  • CHL helium inventory management system operating (Tank Farm, Helium Purifier, A0 GHe Recovery) and Tank Farm has been purified; the helium losses are normal.
  • CHL Nitrogen Reliquefier plant off and LN2 supply dependent on vendor deliveries to CHL LN2 Dewar tanks; deliveries are fine.

FRIG Maintenance

  • 10 out of 13 wet expander flywheels have been replaced
  • 10 of 26 wet/dry expander overhauls of cold ends are completed
  • Various vacuum, instrumentation, and controls maintenance continuing

CHL Maintenance

  • Coldbox-1:
    • Charcoal Adsorber east unit regeneration at high temperature continues with cooldown starting today
    • LHe Subcooler control valve removed for problem diagnosis
  • Nitrogen Reliquefier plant maintenance in progress:
    • Marley Cooling Tower cleanup/overhaul almost finished
    • Rotoflow compressor aftercooler tube bundle being replaced
    • Niagara cooling tower sumps were cleaned
  • Various compressor and facility maintenance in progress

Controls

This Week

  • AD Server Updates complete
  • Database Table Update complete
  • MAC & Computer Room UPS Maintenance complete
  • Acc. Network Switch IOS Updates complete (120)
  • Computer Room Prep work for 8/18 Switch replacement
  • MADC II reset updates continue
  • CAMAC Front End SLD Updates continue
  • Continuing routine maintenance
  • JAVA Engine Restarts/Updates
    • Having to reinstall OS on Solaris workstations

Next Week

  • Saturday 8/18
    • Replace EOL Gigabit Network Switch with Cat6506 Chassis
      • No AD general networking (desktops). 2-4hrs.
  • Tuesday 8/21
    • General Network (desktop) Switch IOS Updates
    • VAX Clusters move to MAC Room
      • 4 clusters, move one at a time, down 30 min/cluster
      • Alarms server, Fileshare, Console page access
  • Wednesday 8/22
    • General Network Switch IOS Updates
    • Upgrade Sup720 to Cat6509
      • No Accelerator Division Network. 2-4hrs.
  • Thursday 8/23
    • General Network Switch IOS Updates
  • Saturday 8/25
    • AD Web Server Updates ~2hr.
      • No Time Server, No Elog access, No node reboot capability

Get Requests In Now

  • Hardware you want to use may not be readily available
    • New infrastructure (links, clocks, networking) takes time
  • Work has to get scheduled

FESS

Casey’s Pond

  • Experts have discovered a leaking 16” valve that needs to be replaced
  • Water will be bypassed through 10” valve during replacement
    • Pressures will drop
    • Experts will perform test with 10” valve before beginning work

Chillers

  • Chiller #3 is off for maintenance
    • HX3 is on as a replacement for Chiller #3
    • There are no back pumps during this work

Poles

  • 12 or 13 of the 19 poles have been replaced

TeV pond silt

  • New intake ports have been built to minimize the silt problem

Safety

  • Safety System Status
    • Upgrade M02 - M05 Enclosure Interlocks 25%
    • Upgrade Main Injector PLC Code and replace PLC Batteries 35%
    • Replace Safety System DEC talk with ATT Natural Voices 15%
    • Linux kernal 2.4 to 2.6 upgrade to Safety Data Acquisition System 25%
    • Upgrade Meson Area Oxygen Monitors 0%
    • Install Radiation Monitor Cables for the Muon Test Area 0%
    • Replace Pbar Vault Key Tree Door Latches & add status LEDs 0%
    • Route Key Tree release signals from MCR to Meson 0%
  • A successful shutdown is getting our work done safely

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 08/24/2007    email Fermilab